how many newspaper does quang delliver in 4 days if he delivers4'200 in three week period[21days]
Answer by checkley71(8403)

4/21=X/4,200 CROSS MULTIPLY
21X=4*4,200
21X=16,800
X=16,9800/21
X=800 DELIVERED IN 4 DAYS.
PROOF
4/21=800/4,200
21*800=4*4,200
16,800=16,800